A battery of emf 60V is connected across a parallel combination of a 8 ohm resistor and a thick wire of negligible resistance.If the current observed is 6A, what is the internal resistance of the battery?

Grade:12th pass

2 Answers

Physics Expert
6 years ago
Calculate the internal resistance of the battery.

Hence, the internal resistance is 12 ohm.
shubham
6 years ago
Internal resistance I=E/R+r
6= 60/8+r
48+6r=60-48 
6r=12 ohm

r=2 ohm
Hence, the internal resistance is 2 ohm.